OpenStudy (anonymous):

On dividing Polynomials 2x^3-9x^2+15/2x-5 thanks in advance

OpenStudy (anonymous):

2x^3-9x^2+15/2x-5 f(x)=2x^3-9x^2+15 by 2x-5 2x-5=0 x=5/2 f(x)=2(2.5)^3-9(2.5)^2+15 f(x)=-10

OpenStudy (anonymous):

is ti right???

OpenStudy (anonymous):

if it is then medal plz

OpenStudy (anonymous):

x^2-2x-5-10/(5-2x)

OpenStudy (akshay_budhkar):

anhhuyalex? what did you do?

OpenStudy (akshay_budhkar):

x^2-2x-5-(10/(5-2x))?

OpenStudy (anonymous):

akshay budhkar, yeah, your way is clearer

OpenStudy (anonymous):

\[2x^3-9x^2+15\div2x-5\]

OpenStudy (akshay_budhkar):

lol i just confirmed? coz i messed it up with the full as numerator and maybe the asker can do the same..

OpenStudy (anonymous):

that's what i originally meant

OpenStudy (akshay_budhkar):

anhhuyalex's answer is right i believe Oglon3r

OpenStudy (anonymous):

x^2-2x+5-(10/2x-5)

OpenStudy (akshay_budhkar):

thats what we r telling lol

OpenStudy (anonymous):

yah thanks!

OpenStudy (anonymous):

lol, you right

OpenStudy (anonymous):

am i right or wrong

OpenStudy (anonymous):

these godamn long divisions......

OpenStudy (akshay_budhkar):

nick i am not getting what you did!

OpenStudy (anonymous):

i did it in polynomial way its in addmaths

OpenStudy (anonymous):

by hoo song thong

OpenStudy (anonymous):

nick got the residual of the division right now that I'm taking a closer look to it

OpenStudy (akshay_budhkar):

WOW! what is that? i heard it for d first time

OpenStudy (anonymous):

thanks everybody

OpenStudy (anonymous):

\[\frac{2x^3-9x^2+15}{2x-5} = x^2 - 2x - 5 - \frac{10}{2x-5}\] Everyone else's answers are wrong.

OpenStudy (anonymous):

http://www.purplemath.com/modules/polys/divani04.gif

OpenStudy (anonymous):

http://www.purplemath.com/modules/polys/div19.gif

OpenStudy (anonymous):

Yes, actually. Akshay answered \[x^2 - 2x - 5 - \frac{10}{5-2x}\] which is most certainly not equal to \[x^2 - 2x - 5 - \frac{10}{2x-5}\] the latter of which is the correct polynomial that results from the division of the two polynomials that you cited in your question.

OpenStudy (anonymous):

on the second gif in the final part they go something like this \[+ -10/2x-5\]

OpenStudy (anonymous):

i answer that too x^2-2x+5-(10/2x-5) look carefully if you want to out the minus you must do carefully + -10/2x-5 if being this - (10/-2x+5) or - (10/5-2x)

OpenStudy (anonymous):

thanks for clarifying it threw me off really bad for some reason haha its been a long night

OpenStudy (anonymous):

no reason to threw you off oglon3r i just explain to lollercakes your explanation is right

OpenStudy (anonymous):

Still wrong, dinainjune. \[-\frac{10}{2x-5} \ne -\frac{10}{-2x+5}\] for all x, which is actually what you're implying.

OpenStudy (akshay_budhkar):

lollercakes i like your confidence and the way you answer! i will be your fan for that.. i am not verifying the answer now though as i am almost asleep but still nice way to answer

OpenStudy (anonymous):

Thanks, Akshay. Dinainjune, I assure you, my answer is correct.

OpenStudy (anonymous):

it's really nice having different answer. okay, if you sure of that lollercakes :)

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=%282x^3-9x^2%2B15%29%2F%282x-5%29 is the answer short cut lol

OpenStudy (anonymous):

And here's the verification of my answer: \[\frac{2x^3 - 9x^2 + 15}{2x - 5} = x^2 - 2x - 5 - \frac{10}{2x-5}\] \[(x^2 - 2x - 5 - \frac{10}{2x - 5})(2x - 5) =\] \[(x^2 + (-2x) + (-5) + (-\frac{10}{2x-5}))(2x - 5) = \] \[x^2(2x - 5) + (-2x)(2x - 5) + (-5)(2x - 5) + (-\frac{10}{2x-5})(2x - 5) = \] \[2x^3 - 5x^2 -4x^2 + 10x - 10x + 25 + (-1)\frac{10}{2x - 5}(2x - 5) = \] \[2x^3 - 9x^2 + 25 + (-1)\frac{10(2x - 5)}{2x - 5} = \] \[2x^3 - 9x^2 + 25 + (-1)*10 = \] \[2x^3 - 9x^2 + 25 - 10 = \] \[2x^3 - 9x^2 + 15\]

OpenStudy (akshay_budhkar):

OMG!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

cool! Thank you for explaining

OpenStudy (anonymous):

Hope it helps. (I'm not trying to be a jerk. I just want everyone to get why the answer is what it is.)

OpenStudy (anonymous):

nope, if you don't do that. i can't find my mistake and repair it Thank you again!